OpenStudy (anonymous):

help solve ( - 1 ) ( - 3 ) - x - 3 7

hero (hero):

I suppose cross cancellation doesn't mean much to you.

hero (hero):

\[-\frac{1}{\cancel{3}} \times -\frac{\cancel{3}}{7}\]

OpenStudy (anonymous):

is the answer -3

hero (hero):

No, I have no idea why you think the answer is -3.

OpenStudy (anonymous):

well i assume its 7

OpenStudy (anonymous):

im sorry \[\frac{ 1 }{ 7 }\]

hero (hero):

Yes, that is correct

OpenStudy (anonymous):

okay could you break it down step by step so i could better understand it please

hero (hero):

Well, the only thing I can do to explain it is this: The general rule for multiplying fractions is: \[\frac{a}{b} \times \frac{c}{d} = \frac{ac}{bd}\] In this case: \[-\frac{1}{3} \times -\frac{3}{7} = \frac{(1)(3)}{(7)(3)} \]

hero (hero):

3/3 cancels leaving just 1/7

hero (hero):

3/3 = 1 so what you really have remaining is \[\frac{1}{7} \times 1\]

hero (hero):

\[\space\space\space\space-\frac{1}{{3}} \times -\frac{{3}}{7} \\=\frac{(1)(3)}{(3)(7)} \\=\frac{(1)(3)}{(7)(3)} \\=\frac{(1)\cancel{(3)}}{(7)\cancel{(3)}} \\=\frac{1}{7} \times 1 \\=\frac{1}{7}\]

hero (hero):

That is about as complete an explanation I can give. If you still don't understand, hire a tutor.

Parth (parthkohli):

Hero would work perfectly as a tutor.